I hope this is posted in the correct place. I need some help I do not understand how to do this. Thanks for any help you can give.
solve by elimination process.
0.3x -0.2y = 4
0.5x + 0.3y = 5
What is the solution of the system. It asks for an ordered pair. Thanks so much.

Answer by stanbon(75887)

solve by elimination process.
0.3x -0.2y = 4
0.5x + 0.3y = 5
-----------------
Multiply thru both equations by 10 to get rid of those decimals:
3x - 2y = 40
5x + 3y = 50
---------------
Multiply the 1st quation by 3 and the 2nd equation by 2 to get:
9x -6y = 120
10x+6y = 100
-----------------
Add the two equations so you can solve for "x":
19x = 220
x = 220/19
--------------
Substitute to solve for Y:
3(220/19) -2y = 40
Multiply thru by 19 to get:
660 -38y = 760
38y = -100
y = -50/19
================
Ordered pair solution: (220/19 , -50/19)
Cheers,
Stan H.